OpenStudy (anonymous):

the number of bus riders form the suburbs to downtown per day is represented by 1200(1.5-x), where x is the fare in dollars. what fare will maximize the total revenue?

OpenStudy (anonymous):

when x is 0, but that's the math. Logically I am not sure. If x = 0 than the equation becomes 1200(1.5) instead of if x = constant than the answer will become smaller.

OpenStudy (anonymous):

r(x)= 1200(1.50-x) 1800-1200x=0 -1200x/-1200=-1800/-1200 x=1.5 is that right?